OpenStudy (anonymous):

In Logs, How do you solve, tan (32°15') ?

OpenStudy (anonymous):

32degree 15 mins = 32+15/60 = 32+1/4 = 32.25 degrees Put that in your calculator and u'l get something around 0.630953

Parth (parthkohli):

For the record, this is Trigonometry, not Logarithms.